# Intrinsic Nonlinear Planar Thermal Hall effect

### v1

@@ Abstract
Introduction— Transport phenomena governed by the quantum geometric properties of Bloch bands have emerged as a central theme in modern condensed matter physics. Over the past decade, concepts such as Berry curvature, orbital moment, and quantum metric have been recognized as key ingredients determining a wide class of Hall-type responses, even in systems without external magnetic fields. These geometrical quantities encapsulate the topological and symmetry characteristics of the underlying crystal, giving rise to effects such as the anomalous Hall effect, anomalous Nernst effect, and nonlinear Hall transport in both magnetic and nonmagnetic materials [1, 2, 3]. These discoveries have established a unified language connecting topology, symmetry, and dissipationless response. These responses arise from two distinct origins: intrinsic mechanisms, rooted in the band geometry itself, and extrinsic mechanisms, which rely on scattering-induced asymmetries of charge carriers [4, 2, 3, 5]. Particularly, the well-known linear anomalous Hall effect is a purely intrinsic phenomenon governed solely by the Berry curvature—a fundamental band geometric quantity—and is independent of the scattering time $\tau$ [3, 6]. In contrast, the second-order anomalous Hall effect, linked to the Berry curvature dipole, explicitly depends on scattering time $\tau$ and is therefore regarded as an extrinsic contribution [7].
@@
Recently, nonlinear Hall phenomena have attracted considerable attention as they extend the concept of Berry-curvature–driven transport beyond the linear-response regime. In particular, the discoveries of the intrinsic second-order anomalous Hall effect [8, 9], the intrinsic second-order valley Hall effect [10], and the extrinsic third-order anomalous Hall effect [11, 12, 13, 14] have revealed that the Berry curvature itself can be field-corrected by an applied electric field through the so called Berry connection polarizability (BCP) tensor or the renormalized quantum metric [15, 16, 17]. Such field-induced corrections expand the established taxonomy of intrinsic Hall responses beyond the linear order. Extending this framework, Huang et al. [18] identified an intrinsic nonlinear planar Hall effect, originating from the BCP. In this effect, the Hall current scales as $j_{H}\propto E^{2}B$, with $E$ representing the driving electric field and $B$ an in-plane magnetic field that couples to the band geometry.
@@
Thermal transport, governed by analogous band geometric principles, offers a natural arena for extending these concepts. It is well established that the anomalous thermal Hall and Nernst effects originate from the Berry curvature weighted by the energy distribution of carriers. However, their nonlinear thermal counterparts remain largely less explored, especially those arising intrinsically from corrections to the quantum geometric structure of Bloch states rather than from extrinsic scattering processes. Analogous to the electric field induced correction of Berry curvature, it has recently been shown that a temperature gradient can also generate a correction to the Berry curvature, characterized by a corresponding thermal Berry connection polarizability (TBCP) [19, 20].
@@
Motivated by this insight, in this work we propose and develop the concept of an intrinsic nonlinear planar thermal Hall effect (NPTHE) – a nonlinear, dissipationless heat current that arises transverse to a temperature gradient as a consequence of a thermal gradient induced correction to the Berry curvature. The planar thermal Hall effect (PTHE) refers to a configuration in which the applied magnetic field ($B$) and temperature gradient ($\nabla T$) both lie within the transport plane that also hosts the transverse heat current $j^{Q}_{H}$. Unlike the conventional thermal Hall effect, where an out-of-plane magnetic field drives a Lorentz-type deflection of carriers, the PTHE occurs entirely under in-plane conditions, where the Lorentz force plays no role. Instead, the intrinsic NPTHE solely originates from the interplay between the in-plane magnetic field and the band geometry, manifesting through a thermal gradient induced correction to the Berry curvature mediated by the thermal Berry connection polarizability (TBCP) [19, 20] and its spin susceptibility [18]. This planar configuration provides a clean platform to probe intrinsic quantum geometric contributions to nonlinear thermal Hall transport, independent of extrinsic scattering effects.
@@
Theoretical formalism and origin of intrinsic NPTHE— To uncover the microscopic origin of the intrinsic NPTHE, we consider a nonmagnetic system subjected to an in-plane temperature gradient $\nabla T$ and a magnetic field $B$ in a coplanar $x$-$y$ configuration. Since the heat current is odd under time reversal $\mathcal{T}$, while $\nabla T$ is even, the intrinsic transport current cannot arise without the presence of the magnetic field. In this coplanar configuration, the magnetic field couples to the electron’s magnetic moment through a Zeeman-like interaction. Consequently, the effect of the magnetic field can be treated as a perturbation to the band structure, inducing spin splitting and effectively converting the original nonmagnetic bands into a “magnetic” band structure. The local Hamiltonian incorporating the Zeeman term can be written as $\hat{H}_{c}=\hat{H}_{0}+\frac{g}{\hbar}\bm{\mu}_{B}\cdot\mathbf{B}$, where $\hat{H}_{0}$ is the unperturbed Hamiltonian, and $g$ and $\mu_{B}$ are the Landé $g$-factor and Bohr magneton respectively. The magnetic field couples to the electron spin, thereby modifying both the Bloch wave function, and the band dispersion, which are designated by $e^{i\mathbf{k\cdot r}}|\widetilde{u}_{n\bm{k}}\rangle$, and $\widetilde{\varepsilon}_{nk}=\langle\widetilde{u}_{nk}|[\hat{H}_{0}+\frac{g}{\hbar}\bm{\mu}_{B}\cdot\mathbf{B}]|\widetilde{u}_{nk}\rangle$, respectively. Here, $n$ is the quantum number representing band index. In this Bloch basis, the interband and intraband Berry connections can be defined as $\bm{\mathcal{\tilde{A}}}_{mn}=\langle\widetilde{u}_{m\bm{k}}|i\partial_{\bm{k}}|\widetilde{u}_{n\bm{k}}\rangle$ and $\bm{\mathcal{\tilde{A}}}_{n}=\langle\widetilde{u}_{n\bm{k}}|i\partial_{\bm{k}}|\widetilde{u}_{n\bm{k}}\rangle$, respectively, with the corresponding Berry curvature given by $\widetilde{\bm{\Omega}}_{n\bm{k}}=\bm{\nabla}_{\bm{k}}\times\bm{\mathcal{\tilde{A}}}_{n\bm{k}}$. Moreover, the presence of a temperature gradient $\nabla T$ in the thermal Hall setup introduces an additional perturbation term $\hat{H}^{{}^{\prime}}=-\frac{1}{2}\{\hat{H}_{c},\hat{r}\}\cdot\bm{E}^{T}$ to the local Hamiltonian $\hat{H}_{c}$, where $\hat{r}$ is the position operator and $\bm{E}^{T}=-\frac{\nabla T}{T}$ represents the thermal field [19, 20]. Using time-independent perturbation theory, the first-order correction to the wave function $|\widetilde{u}_{n\bm{k}}\rangle$ due to $\bm{E}^{T}$ field gives rise to a thermal field induced modification of the Berry curvature, $\delta^{T}\widetilde{\bm{\Omega}}_{n\bm{k}}=\nabla_{\bm{k}}\times\bm{\mathcal{\tilde{A}}}_{n\bm{k}}^{T}$ and hence the total Berry curvature upto first-order in $\bm{E}^{T}$ field is given by $(\widetilde{\bm{\Omega}}_{n\bm{k}}+\delta^{T}\widetilde{\bm{\Omega}}_{n\bm{k}})$. Consequently, the Bloch state $|\widetilde{u}_{n\bm{k}}\rangle$ also acquires a second-order energy correction $\delta^{T}\widetilde{\varepsilon}_{n\bm{k}}=-\frac{1}{2}\bm{\mathcal{\tilde{A}}}_{n}^{T}\cdot\bm{E}^{T}$. Here, $\bm{\mathcal{\tilde{A}}}_{n}^{T}$ denotes the first-order correction to the Berry connection due to thermal field. $\bm{\mathcal{\tilde{A}}}_{n}^{T}$ depends on the quantum geometry of the Bloch function, and is expressed as $\bm{\mathcal{\tilde{A}}}_{n,a}^{T}=\widetilde{\mathcal{F}}_{ab}^{n}\bm{E}^{T}_{b}$, where $\widetilde{\mathcal{F}}_{ab}^{n}$ is a gauge invariant quantity known as the thermal Berry curvature polarizability (TBCP) tensor [19, 20, 21] associated with the state $|\widetilde{u}_{n\bm{k}}\rangle$. The indices a,ba,\,b represent the cartesian components and the repeated indices are summed over. The detailed derivations of these expressions are provided in the Supplementary Material (SM) [21]. Since the correction to the Berry curvature, $\delta^{T}\widetilde{\bm{\Omega}}_{n\bm{k}}$ is already linear in the thermal gradient $\nabla T$, the corresponding heat current generated by this term naturally appears at second order in $\nabla T$. The resulting nonlinear intrinsic thermal Hall current takes the form [21, 22, 23, 24],
@@
where $f_{0}(\widetilde{\varepsilon}_{n\bm{k}})$ is the Fermi–Dirac distribution, $\mu$ the chemical potential, $\mathrm{k}_{B}$ the Boltzmann constant, and  $\text{ Li}_{2}$ the dilogarithm function. This expression captures the intrinsic nonlinear planar thermal Hall current, arising purely from the thermally corrected Berry curvature and independent of scattering processes. The corresponding second-order thermal Hall response tensor is defined as $j^{Q(2,\,\text{int})}_{a}=\kappa_{abc}\left(-\nabla T_{b}\right)\left(-\nabla T_{c}\right)$, where $\kappa_{abc}$ encodes the nonlinear thermal conductivity components. In Eq. (1), the material-specific information is embedded in $\widetilde{\mathcal{F}}_{ab}^{n}$ and the energy dispersion $\widetilde{\varepsilon}_{n\bm{k}}$, both defined with respect to the magnetic-field–perturbed band structure and Bloch state $|\widetilde{u}_{n\bm{k}}\rangle$. By straightforward expansion (see SM [21]), these quantities can be expressed perturbatively in powers of the magnetic field $B$, allowing them to be related to the original unperturbed Bloch state $|u_{n\bm{k}}\rangle$ and energy eigenvalue $\varepsilon_{n\bm{k}}$. To first order in $B$, we obtain
@@
where $\mathcal{F}_{ab}^{n}$ is the thermal Berry connection polarizability (TBCP) of the $n^{th}$ band in the basis of unperturbed Bloch state $|u_{n\bm{k}}\rangle$, and $\Lambda_{abc}^{n}=\partial_{B_{c}}\widetilde{\mathcal{F}}_{ab}^{n}\big|_{B=0}$ can be interpreted as the spin susceptibility of the TBCP, evaluated for the unperturbed Bloch state. The explicit expressions for $\mathcal{F}_{ab}^{n}$ and $\Lambda_{abc}^{n}$ are given in Eq. (3) (we set $e=\hbar=1$), while the detailed derivations are presented in the SM [21].
@@
Here, the indices belong to cartesian components $\{x,y\}$ and $d\bm{k}]$ represents the two-dimensional Brillouin-zone integral $d^{2}k/(2\pi)^{2}$. The detailed derivations of the above expressions are provided in the SM [21]. Evidently, the conductivity tensor $\kappa_{abcd}^{\text{int}}$ is independent of scattering time $\tau$; hence, as discussed earlier, it constitutes an intrinsic contribution to the nonlinear thermal conductivity. This intrinsic response is distinct from extrinsic mechanisms and is fundamentally different in origin from the previously reported nonlinear electrical and thermal Hall conductivities [25, 26, 27, 28].
@@
where $s=\pm$ is the valley index, $t_{s}=st$ parametrizes the valley dependent tilt, $v_{x,y}$ are the Fermi velocities along the principal axes, $m$ is the Dirac mass that opens an energy gap, $\sigma_{i}$’s are the Pauli matrices, and $\sigma_{0}$ is the identity matrix. This minimal model captures the low-energy physics of a broad class of two-dimensional Dirac materials [32, 33], including the surfaces of topological crystalline insulators [34, 35, 36, 37] and strained transition-metal dichalcogenides [38, 39, 40]. The inclusion of the tilt parameter $t$ in Eq. (8) breaks both rotational and inversion symmetries. When both valleys are considered together, the full Hamiltonian retains a mirror symmetry $M_{x}:(k_{x}\rightarrow-k_{x})$, which interchanges the two valleys. For vanishing mass term each individual valley separately preserves an additional mirror symmetry $M_{y}$. These mirror symmetries play crucial roles in determining the allowed components of the intrinsic NPTHE as discussed in Table 1. The energy dispersion of the Hamiltonian is given by $\varepsilon_{k}^{\pm}=stk_{x}\pm(m^{2}+k_{x}^{2}v_{x}^{2}+k_{y}^{2}v_{y}^{2}),^{1/2}$ where $\pm$ denotes the conduction and valence band, respectively. The TBCP tensor for this Dirac Hamiltonian can be directly evaluated using Eq. (3). Interestingly, the TBCP is proportional to the valley-dependent tilt parameter $t_{s}$, in contrast to the electric field induced BCP [14, 18, 10]. The explicit analytical expressions for the components of the TBCP tensors are provided in the SM [21]. Note that, owing to the mirror symmetry $M_{x}$, which interchanges the two valleys, the intrinsic NPTHE tensor satisfies $\kappa_{xyyy}^{\text{int}}(s,k_{x},k_{y})=-\kappa_{xyyy}^{\text{int}}(-s,-k_{x},k_{y})$ as can be directly verified from the analytical structure of the integrand in Eq. (5). This relation implies that the valley-resolved contributions of $\kappa_{xyyy}^{\text{int}}$ are opposite in sign, i.e., $\kappa_{xyyy}^{\text{int}}(s)=-\kappa_{xyyy}^{\text{int}}(-s)$ due to the presence of $M_{x}$ mirror. In contrast, the component $\kappa_{yxxx}^{\text{int}}$ remains even under valley exchange, satisfying $\kappa_{yxxx}^{\text{int}}(s)=\kappa_{yxxx}^{\text{int}}(-s)$. Furthermore, the integrand of $\kappa_{abba}^{\text{int}}(s,k_{x},k_{y})$ is odd in $k_{y}$, and therefore its contribution vanishes upon Brillouin-zone integration. Focusing on the $s=+$ valley, the analytical expressions for the components of $\kappa_{abbb}^{\text{int}}$ in the conduction band (chemical potential $\mu>0$) for the isotropic case $v_{x}=v_{y}=v$ are obtained as
@@
Outlook— The intrinsic NPTHE proposed here can be experimentally distinguished from other nonlinear Hall responses–such as the Berry curvature dipole driven nonlinear Hall effect [24, 41] or extrinsic skew-scattering [42, 13] contributions–through its distinct scaling behavior, symmetry dependence, and dissipationless nature. First, unlike extrinsic mechanisms that depend on the relaxation time $\tau$ [24, 41, 42, 43, 44, 25], the intrinsic NPTHE originates purely from the quantum geometric correction to the Berry curvature and is therefore independent of $\tau$, similar to the intrinsic anomalous Hall and thermal Hall effects [3, 1]. This distinction can be verified experimentally by examining its temperature and disorder dependence: a $\tau$-independent response persisting in the clean limit signals intrinsic origin. Second, the planar configuration–with the magnetic field and temperature gradient lying in the same plane–provides a clear diagnostic distinction from ordinary thermal Hall effects, where the field is perpendicular to the transport plane. The angular modulation of the Hall voltage with respect to the relative orientation of the thermal and magnetic fields, following the $\cos(\phi-\theta)$ or $\sin(\phi-\theta)$ dependence predicted in Eq. (7), would serve as a direct signature of the planar geometry. Furthermore, the symmetry selection rules derived earlier imply that the signal should vanish in centrosymmetric or mirror-symmetric crystals (e.g., $C_{3h}$, $D_{3h}$), but appear exclusively in polar or chiral point groups lacking horizontal mirror symmetry–providing an additional symmetry-based criterion for identification. Together, these features offer experimentally accessible means to isolate and confirm the intrinsic NPTHE as a distinct geometric nonlinear transport phenomenon.
@@
I am grateful to Dr. Snehasish Nandy, Dr. Surajit Sarkar, Dr. Hridis K Pal, and Bishal Das for valuable discussions. I acknowledge the Department of Physics at University of Cagliari, Italy and Dr. Fabio Bernardini, & Dr. Alessio Filippetti for providing the necessary facilities.

### v2

@@ Abstract
Introduction— Transport phenomena governed by the quantum geometric properties of Bloch bands have emerged as a central theme in modern condensed matter physics. Over the past decade, concepts such as Berry curvature, orbital moment, and quantum metric have been recognized as key ingredients determining a wide class of Hall-type responses, even in systems without external magnetic fields. These geometrical quantities encapsulate the topological and symmetry characteristics of the underlying crystal, giving rise to effects such as the anomalous Hall effect, anomalous Nernst effect, and nonlinear Hall transport in both magnetic and nonmagnetic materials [Xiao2010Berry, Nonlinear2021Du, Nagaosa2010Anomalous]. These discoveries have established a unified language connecting topology, symmetry, and dissipationless response. These responses arise from two distinct origins: intrinsic mechanisms, rooted in the band geometry itself, and extrinsic mechanisms, which rely on scattering-induced asymmetries of charge carriers [sinitsyn2007semiclassical, Nonlinear2021Du, Nagaosa2010Anomalous, Quantum2021Du]. Particularly, the well-known linear anomalous Hall effect is a purely intrinsic phenomenon governed solely by the Berry curvature—a fundamental band geometric quantity—and is independent of the scattering time $\tau$ [Nagaosa2010Anomalous, Jungwirth2002AHE]. In contrast, the second-order anomalous Hall effect, linked to the Berry curvature dipole, explicitly depends on scattering time $\tau$ and is therefore regarded as an extrinsic contribution [Sodemann2015BCD].
@@
Recently, nonlinear Hall phenomena have attracted considerable attention as they extend the concept of Berry-curvature–driven transport beyond the linear-response regime. In particular, the discoveries of the intrinsic second-order anomalous Hall effect [liu2021Intrinsic, wang2021Intrinsic], the intrinsic second-order valley Hall effect [kamaldas2024valleyHall], and the extrinsic third-order anomalous Hall effect [lai2021third, tanay2023thirdorder, barman2025thirdorder, BCP2022Liu] have revealed that the Berry curvature itself can be field-corrected by an applied electric field through the so called Berry connection polarizability (BCP) tensor or the renormalized quantum metric [Gao2014Field, yu2025quantum, gao2023quantum]. Such field-induced corrections expand the established taxonomy of intrinsic Hall responses beyond the linear order. Extending this framework, Huang et al. [IPHE2023] identified an intrinsic nonlinear planar Hall effect, originating from the BCP. In this effect, the Hall current scales as $j_{H}\propto E^{2}B$, with $E$ representing the driving electric field and $B$ an in-plane magnetic field that couples to the band geometry.
@@
Thermal transport, governed by analogous band geometric principles, offers a natural arena for extending these concepts. It is well established that the anomalous thermal Hall and Nernst effects originate from the Berry curvature weighted by the energy distribution of carriers. However, their nonlinear thermal counterparts remain largely less explored, especially those arising intrinsically from corrections to the quantum geometric structure of Bloch states rather than from extrinsic scattering processes. Analogous to the electric field induced correction of Berry curvature, it has recently been shown that a temperature gradient can also generate a correction to the Berry curvature, characterized by a corresponding thermal Berry connection polarizability (TBCP) [li2024TBCP, zhang2025TBCP].
@@
Motivated by this insight, in this work we propose and develop the concept of an intrinsic nonlinear planar thermal Hall effect (NPTHE) – a nonlinear, dissipationless heat current that arises transverse to a temperature gradient as a consequence of a thermal gradient induced correction to the Berry curvature. The planar thermal Hall effect (PTHE) refers to a configuration in which the applied magnetic field ($B$) and temperature gradient ($\nabla T$) both lie within the transport plane that also hosts the transverse heat current $j^{Q}_{H}$. Unlike the conventional thermal Hall effect, where an out-of-plane magnetic field drives a Lorentz-type deflection of carriers, the PTHE occurs entirely under in-plane conditions, where the Lorentz force plays no role. Instead, the intrinsic NPTHE solely originates from the interplay between the in-plane magnetic field and the band geometry, manifesting through a thermal gradient induced correction to the Berry curvature mediated by the thermal Berry connection polarizability (TBCP) [li2024TBCP, zhang2025TBCP] and its spin susceptibility [IPHE2023]. This planar configuration provides a clean platform to probe intrinsic quantum geometric contributions to nonlinear thermal Hall transport, independent of extrinsic scattering effects.
@@
Theoretical formalism and origin of intrinsic NPTHE— To uncover the microscopic origin of the intrinsic NPTHE, we consider a nonmagnetic system subjected to an in-plane temperature gradient $\nabla T$ and a magnetic field $B$ in a coplanar $x$-$y$ configuration. Since the heat current is odd under time reversal $\mathcal{T}$, while $\nabla T$ is even, the intrinsic transport current cannot arise without the presence of the magnetic field. In this coplanar configuration, the magnetic field couples to the electron’s magnetic moment through a Zeeman-like interaction. Consequently, the effect of the magnetic field can be treated as a perturbation to the band structure, inducing spin splitting and effectively converting the original nonmagnetic bands into a “magnetic” band structure. The local Hamiltonian incorporating the Zeeman term can be written as $\hat{H}_{c}=\hat{H}_{0}+\frac{g}{\hbar}\bm{\mu}_{B}\cdot\mathbf{B}$, where $\hat{H}_{0}$ is the unperturbed Hamiltonian, and $g$ and $\mu_{B}$ are the Landé $g$-factor and Bohr magneton respectively. The magnetic field couples to the electron spin, thereby modifying both the Bloch wave function, and the band dispersion, which are designated by $e^{i\mathbf{k\cdot r}}|\widetilde{u}_{n\bm{k}}\rangle$, and $\widetilde{\varepsilon}_{nk}=\langle\widetilde{u}_{nk}|[\hat{H}_{0}+\frac{g}{\hbar}\bm{\mu}_{B}\cdot\mathbf{B}]|\widetilde{u}_{nk}\rangle$, respectively. Here, $n$ is the quantum number representing band index. In this Bloch basis, the interband and intraband Berry connections can be defined as $\bm{\mathcal{\tilde{A}}}_{mn}=\langle\widetilde{u}_{m\bm{k}}|i\partial_{\bm{k}}|\widetilde{u}_{n\bm{k}}\rangle$ and $\bm{\mathcal{\tilde{A}}}_{n}=\langle\widetilde{u}_{n\bm{k}}|i\partial_{\bm{k}}|\widetilde{u}_{n\bm{k}}\rangle$, respectively, with the corresponding Berry curvature given by $\widetilde{\bm{\Omega}}_{n\bm{k}}=\bm{\nabla}_{\bm{k}}\times\bm{\mathcal{\tilde{A}}}_{n\bm{k}}$. Moreover, the presence of a temperature gradient $\nabla T$ in the thermal Hall setup introduces an additional perturbation term $\hat{H}^{{}^{\prime}}=-\frac{1}{2}\{\hat{H}_{c},\hat{r}\}\cdot\bm{E}^{T}$ to the local Hamiltonian $\hat{H}_{c}$, where $\hat{r}$ is the position operator and $\bm{E}^{T}=-\frac{\nabla T}{T}$ represents the thermal field [li2024TBCP, zhang2025TBCP]. Using time-independent perturbation theory, the first-order correction to the wave function $|\widetilde{u}_{n\bm{k}}\rangle$ due to $\bm{E}^{T}$ field gives rise to a thermal field induced modification of the Berry curvature, $\delta^{T}\widetilde{\bm{\Omega}}_{n\bm{k}}=\nabla_{\bm{k}}\times\bm{\mathcal{\tilde{A}}}_{n\bm{k}}^{T}$ and hence the total Berry curvature upto first-order in $\bm{E}^{T}$ field is given by $(\widetilde{\bm{\Omega}}_{n\bm{k}}+\delta^{T}\widetilde{\bm{\Omega}}_{n\bm{k}})$. Consequently, the Bloch state $|\widetilde{u}_{n\bm{k}}\rangle$ also acquires a second-order energy correction $\delta^{T}\widetilde{\varepsilon}_{n\bm{k}}=-\frac{1}{2}\bm{\mathcal{\tilde{A}}}_{n}^{T}\cdot\bm{E}^{T}$. Here, $\bm{\mathcal{\tilde{A}}}_{n}^{T}$ denotes the first-order correction to the Berry connection due to thermal field. $\bm{\mathcal{\tilde{A}}}_{n}^{T}$ depends on the quantum geometry of the Bloch function, and is expressed as $\bm{\mathcal{\tilde{A}}}_{n,a}^{T}=\widetilde{\mathcal{F}}_{ab}^{n}\bm{E}^{T}_{b}$, where $\widetilde{\mathcal{F}}_{ab}^{n}$ is a gauge invariant quantity known as the thermal Berry curvature polarizability (TBCP) tensor [li2024TBCP, zhang2025TBCP, supp] associated with the state $|\widetilde{u}_{n\bm{k}}\rangle$. The indices a,ba,\,b represent the cartesian components and the repeated indices are summed over. The detailed derivations of these expressions are provided in the Supplementary Material (SM) [supp]. Since the correction to the Berry curvature, $\delta^{T}\widetilde{\bm{\Omega}}_{n\bm{k}}$ is already linear in the thermal gradient $\nabla T$, the corresponding heat current generated by this term naturally appears at second order in $\nabla T$. The resulting nonlinear intrinsic thermal Hall current takes the form [supp, HeatCurrent2010Bergman, HeatCurrent2011Yokoyama, NATHE2020Nandy],
@@
where $f_{0}(\widetilde{\varepsilon}_{n\bm{k}})$ is the Fermi–Dirac distribution, $\mu$ the chemical potential, $\mathrm{k}_{B}$ the Boltzmann constant, and  $\text{ Li}_{2}$ the dilogarithm function. This expression captures the intrinsic nonlinear planar thermal Hall current, arising purely from the thermally corrected Berry curvature and independent of scattering processes. The corresponding second-order thermal Hall response tensor is defined as $j^{Q(2,\,\text{int})}_{a}=\kappa_{abc}\left(-\nabla T_{b}\right)\left(-\nabla T_{c}\right)$, where $\kappa_{abc}$ encodes the nonlinear thermal conductivity components. In Eq. (1), the material-specific information is embedded in $\widetilde{\mathcal{F}}_{ab}^{n}$ and the energy dispersion $\widetilde{\varepsilon}_{n\bm{k}}$, both defined with respect to the magnetic-field–perturbed band structure and Bloch state $|\widetilde{u}_{n\bm{k}}\rangle$. By straightforward expansion (see SM [supp]), these quantities can be expressed perturbatively in powers of the magnetic field $B$, allowing them to be related to the original unperturbed Bloch state $|u_{n\bm{k}}\rangle$ and energy eigenvalue $\varepsilon_{n\bm{k}}$. To first order in $B$, we obtain
@@
where $\mathcal{F}_{ab}^{n}$ is the thermal Berry connection polarizability (TBCP) of the $n^{th}$ band in the basis of unperturbed Bloch state $|u_{n\bm{k}}\rangle$, and $\Lambda_{abc}^{n}=\partial_{B_{c}}\widetilde{\mathcal{F}}_{ab}^{n}\big|_{B=0}$ can be interpreted as the spin susceptibility of the TBCP, evaluated for the unperturbed Bloch state. The explicit expressions for $\mathcal{F}_{ab}^{n}$ and $\Lambda_{abc}^{n}$ are given in Eq. (3) (we set $e=\hbar=1$), while the detailed derivations are presented in the SM [supp].
@@
Here, the indices belong to cartesian components $\{x,y\}$ and $d\bm{k}]$ represents the two-dimensional Brillouin-zone integral $d^{2}k/(2\pi)^{2}$. The detailed derivations of the above expressions are provided in the SM [supp]. Evidently, the conductivity tensor $\kappa_{abcd}^{\text{int}}$ is independent of scattering time $\tau$; hence, as discussed earlier, it constitutes an intrinsic contribution to the nonlinear thermal conductivity. This intrinsic response is distinct from extrinsic mechanisms and is fundamentally different in origin from the previously reported nonlinear electrical and thermal Hall conductivities [NPHE2019Pan, NPHE2021Wen, PHE2021Battilomo, PHE2019Nandy].
@@
where $s=\pm$ is the valley index, $t_{s}=st$ parametrizes the valley dependent tilt, $v_{x,y}$ are the Fermi velocities along the principal axes, $m$ is the Dirac mass that opens an energy gap, $\sigma_{i}$’s are the Pauli matrices, and $\sigma_{0}$ is the identity matrix. This minimal model captures the low-energy physics of a broad class of two-dimensional Dirac materials [RevModPhys2010ZahidHasan, RevModPhys2011Xiao], including the surfaces of topological crystalline insulators [TCI2013LiangFu, TCI2014LiangFu, TCI2013Okada, TCI2012Hsieh] and strained transition-metal dichalcogenides [WTe2018Du, TMDC2019Zhou, TMDC2020Zhou]. The inclusion of the tilt parameter $t$ in Eq. (8) breaks both rotational and inversion symmetries. When both valleys are considered together, the full Hamiltonian retains a mirror symmetry $M_{x}:(k_{x}\rightarrow-k_{x})$, which interchanges the two valleys. For vanishing mass term each individual valley separately preserves an additional mirror symmetry $M_{y}$. These mirror symmetries play crucial roles in determining the allowed components of the intrinsic NPTHE as discussed in Table 1. The energy dispersion of the Hamiltonian is given by $\varepsilon_{k}^{\pm}=stk_{x}\pm(m^{2}+k_{x}^{2}v_{x}^{2}+k_{y}^{2}v_{y}^{2}),^{1/2}$ where $\pm$ denotes the conduction and valence band, respectively. The TBCP tensor for this Dirac Hamiltonian can be directly evaluated using Eq. (3). Interestingly, the TBCP is proportional to the valley-dependent tilt parameter $t_{s}$, in contrast to the electric field induced BCP [BCP2022Liu, IPHE2023, kamaldas2024valleyHall]. The explicit analytical expressions for the components of the TBCP tensors are provided in the SM [supp]. Note that, owing to the mirror symmetry $M_{x}$, which interchanges the two valleys, the intrinsic NPTHE tensor satisfies $\kappa_{xyyy}^{\text{int}}(s,k_{x},k_{y})=-\kappa_{xyyy}^{\text{int}}(-s,-k_{x},k_{y})$ as can be directly verified from the analytical structure of the integrand in Eq. (5). This relation implies that the valley-resolved contributions of $\kappa_{xyyy}^{\text{int}}$ are opposite in sign, i.e., $\kappa_{xyyy}^{\text{int}}(s)=-\kappa_{xyyy}^{\text{int}}(-s)$ due to the presence of $M_{x}$ mirror. In contrast, the component $\kappa_{yxxx}^{\text{int}}$ remains even under valley exchange, satisfying $\kappa_{yxxx}^{\text{int}}(s)=\kappa_{yxxx}^{\text{int}}(-s)$. Furthermore, the integrand of $\kappa_{abba}^{\text{int}}(s,k_{x},k_{y})$ is odd in $k_{y}$, and therefore its contribution vanishes upon Brillouin-zone integration. Focusing on the $s=+$ valley, the analytical expressions for the components of $\kappa_{abbb}^{\text{int}}$ in the conduction band (chemical potential $\mu>0$) for the isotropic case $v_{x}=v_{y}=v$ are obtained as
@@
Outlook— The intrinsic NPTHE proposed here can be experimentally distinguished from other nonlinear Hall responses–such as the Berry curvature dipole driven nonlinear Hall effect [NATHE2020Nandy, NTHE2022Nandy] or extrinsic skew-scattering [NTHE2022Zhou, barman2025thirdorder] contributions–through its distinct scaling behavior, symmetry dependence, and dissipationless nature. First, unlike extrinsic mechanisms that depend on the relaxation time $\tau$ [NATHE2020Nandy, NTHE2022Nandy, NTHE2022Zhou, INTHE2023Harsh, LNPHE2021Wen, NPHE2019Pan], the intrinsic NPTHE originates purely from the quantum geometric correction to the Berry curvature and is therefore independent of $\tau$, similar to the intrinsic anomalous Hall and thermal Hall effects [Nagaosa2010Anomalous, Xiao2010Berry]. This distinction can be verified experimentally by examining its temperature and disorder dependence: a $\tau$-independent response persisting in the clean limit signals intrinsic origin. Second, the planar configuration–with the magnetic field and temperature gradient lying in the same plane–provides a clear diagnostic distinction from ordinary thermal Hall effects, where the field is perpendicular to the transport plane. The angular modulation of the Hall voltage with respect to the relative orientation of the thermal and magnetic fields, following the $\cos(\phi-\theta)$ or $\sin(\phi-\theta)$ dependence predicted in Eq. (7), would serve as a direct signature of the planar geometry. Furthermore, the symmetry selection rules derived earlier imply that the signal should vanish in centrosymmetric or mirror-symmetric crystals (e.g., $C_{3h}$, $D_{3h}$), but appear exclusively in polar or chiral point groups lacking horizontal mirror symmetry–providing an additional symmetry-based criterion for identification. Together, these features offer experimentally accessible means to isolate and confirm the intrinsic NPTHE as a distinct geometric nonlinear transport phenomenon.
@@
I am grateful to Dr. Snehasish Nandy, Dr. Surajit Sarkar, Dr. Hridis K Pal, and Bishal Das for valuable discussions. I acknowledge the Department of Physics at University of Cagliari, Italy and Dr. Fabio Bernardini, & Dr. Alessio Filippetti for providing the necessary facilities. The author acknowledges support from the Italian Ministry of University and Research (MUR), financed by the European Union – Next Generation EU, through the PRIN 2022 project SUBLI “Sustainable spin generators based on Van der Waals dichalcogenides,” contract No. 2022M3WXE7. Additional support from the PRIN 2022 TOTEM project, grant No. F53D23001080006, funded by MUR, is also acknowledged.